# Global Control for Local SO(3)-Equivariant Scale-Invariant Vessel Segmentation

### v1

@@ 3.3 Modifying boundary conditions
The global controller we present provides flexible incorporation of custom sections of vessels to include in the 3D vascular model. To demonstrate this, we extract 3D models while adapting $\Omega_{\text{aorta}}$. We include three different regions in $\Omega_{aorta}$ to terminate iterative segmentation in the superior direction: above the top of the T10 vertebra, above the top of the T12 vertebra, and the heart. Figure 4 shows visualizations of the resulting 3D vascular models. Although the iterative tracker was trained only on contours of the abdominal aorta between the T12 vertebra and iliac bifurcation, it still tracks and segments the aortic arch due to its symmetry-preserving properties.
@@ 4 Conclusion & Discussion
We have presented a general framework for automatically extracting smooth, topologically correct vascular models, combining strengths of global and local vessel segmentation methods. A global controller provides boundary conditions for a local iterative vessel segmentation model, that exploits local scale- and rotational symmetries to generalise to vessels of unseen sizes and orientations. Our post-processing step provides a neural field of the vascular model that can be reconstructed to a mesh of any desired resolution. We demonstrated the effectiveness of symmetry preservations by automatic extraction of vascular models of abdominal aortic aneurysms, iliac arteries and renal arteries.
By leveraging local scale- and rotational symmetries, our iterative joint tracking and segmentation module generalizes to vessels of sizes (Fig. 3, left) and sections (Fig. 4) unseen during training. Despite this strong generalisation, some symmetries cannot be preserved. For example, the contour regression model trained on renal arteries struggles to segment the abdominal aorta lumen if surrounded by a thrombus, as this was not seen during training. Moreover, generalisation to other image modalities, e.g. MRI or ultrasound requires retraining. However, as the local vessel segmentation step can be trained on contour annotations of a few patients, data collection and annotation costs are low.
The global controller we present in this work gives the user control over which vessel sections to incorporate in the final vascular model. In combination with the generalising properties of the iterative tracker, this allows to automatically build 3D vascular models in other anatomical regions by providing a different set of heuristics. We used segmentation masks from TotalSegmentator [23] as input for our global controller, an off-the-shelf tool. In combination with the generalising local segmentation method, it has the potential to automatically segment any vascular region without the need for additional specific annotations. Nevertheless, another model could be added to provide more specialized boundary conditions for tracking, e.g. for $p_{\text{renal}}$ in our experiments. Moreover, as all vessel segments are tracked separately using their own set of predefined boundary conditions, the vascular model is fully labelled. As the boundary conditions for each artery are standardized, the regions of interest of the resulting 3D vascular models are consistent across subjects. This enables large-scale studies of the role of hemodynamic and morphological features in the prognosis of cardiovascular diseases.

### v2

@@ 3.3 Modifying boundary conditions
The global controller we present provides flexible inclusion of different vessel ROIs in the 3D model. To show this, we extract multiple 3D models while adapting $\Omega_{\text{aorta}}$. We include three different regions in $\Omega_{aorta}$ to terminate iterative segmentation in the superior direction: until the T10 vertebra, the T12 vertebra, and the heart. Figure 4 shows visualizations of the resulting 3D vascular models. Although the iterative tracker was trained only on contours of the aorta between the T12 vertebra and iliac bifurcation, it still accurately tracks and segments the aortic arch due to its symmetry-preserving properties.
@@ 4 Conclusion & Discussion
We have presented a general framework for automatically extracting smooth, topologically correct vascular models, combining strengths of global and local vessel segmentation methods. A global controller provides boundary conditions for a local iterative vessel segmentation model, that exploits local scale- and rotational symmetries to generalise to vessels of unseen sizes and orientations. Our post-processing step provides a neural field of the vascular model that can be reconstructed to a watertight mesh of any desired resolution. We demonstrated the effectiveness of symmetry preservations by automatic extraction of vascular models of abdominal aortic aneurysms, iliac arteries and renal arteries.
By leveraging local scale- and rotational symmetries, our iterative joint tracking and segmentation module generalizes to vessels of sizes (Fig. 3, left) and segments (Fig. 4) unseen during training. Despite this strong generalisation, some symmetries cannot be preserved. For example, the contour regression model trained on renal arteries struggles to segment the abdominal aorta lumen if surrounded by a thrombus, as this was not seen during training. Moreover, generalisation to other image modalities, e.g. MRI or ultrasound requires retraining. However, as the local vessel segmentation step can be trained on contour annotations of a few patients, data collection and annotation costs are low.
The global controller we present in this work gives control over vessel ROIs to incorporate in the final 3D model. In combination with the generalising properties of the local iterative tracker, this allows us to automatically build 3D vascular models in other anatomical regions by providing a different set of heuristics. We used coarse segmentation masks from TotalSegmentator [24] as input for our global controller, an off-the-shelf tool. In combination with the generalising local segmentation method, it has the potential to automatically segment any vascular region without the need for additional specific annotations. A potential drawback of this approach is that the local segmentation model can only segment those vessel segments for which at least a seed point can be identified. We found, that in some cases, the global controller could not identify renal arteries, thereby affecting overall segmentation performance. Note that we use TotalSegmentator but are not dependent on this specific model. Any existing model that provides binary segmentation masks or seed points could be used, and multiple models could be combined for specialized tasks.
As the boundary conditions for each artery are standardized, the ROIs of the resulting 3D vascular models are consistent across subjects. This enables large-scale studies of the role of hemodynamic and morphological features in the prognosis of cardiovascular diseases and could in future work be extended to quantification of other vessel-specific structures such as intraluminal thrombus and plaques that are important in the analysis of cardiovascular diseases.